# Diffraction Aided Wireless Positioning

### v1

@@ VI Conclusion
In this paper, we investigate positioning in NLOS scenarios and investigate diffraction as a signal propagation mechanism. Using the Geometrical Theory of Diffraction, we develop a new path model for NLOS scenarios where a propagation path is generated due to interaction with an edge. We create a positioning technique based on this approach that can inherently estimate the position of the diffraction point on the diffracting edge. Using Fisher information analysis, we present the necessary and sufficient conditions to estimate the 3D position. Further, we calculate the CRLB on the 3D positioning performance. This positioning approach is then applied to a practical scenario relevant to public safety where there is a need to improve both 3D and Z-axis position estimates. For this, we use a realistic building model that provides insights into signal propagation in the O2I scenario. We present system insights where we control the power in the MPCs generated due to interactions in the window using electric field polarization at the transmitter. Next, we illustrate various ways to achieve the necessary and sufficient conditions to enable 3D position estimates for this system. Finally, we present some numerical results on the lower bound of the positioning performance.
@@
Consider Fig. 1(a); our goal is to derive the geometrical path length $AQ_{e}B$. We start by calculating the coordinates of the diffraction point $Q_{e}$. For this, we substitute the incident ray unit vector $\hat{s^{\prime}}$, edge unit vector $\hat{e}$ and diffracted ray unit vector $\hat{s}$ in terms of the various coordinates shown in Fig. 1(a) in the diffraction law eq. (2). Next, we substitute the parametric coordinates of diffraction point $Q_{e}$ in terms of the parameter $\lambda$ and the endpoints of the edge $X_{1}$ and $X_{2}$. After some algebraic manipulations, we obtain a quadratic equation for the parameter $\lambda$. All these steps are shown in eq. (21).
$\begin{split}&\left|\frac{(q_{x}-x_{a})\hat{x}-y_{a}\hat{y}+(z_{e}-z_{a})\hat{z}}{\sqrt{(q_{x}-x_{a})^{2}+y_{a}^{2}+(z_{e}-z_{a})^{2}}}\cdot\hat{x}\right|=\left|\frac{(x_{n}-q_{x})\hat{x}+y_{n}\hat{y}+(z_{n}-z_{e})\hat{z}}{\sqrt{(x_{n}-q_{x})^{2}+y_{n}^{2}+(z_{n}-z_{e})^{2}}}\cdot\hat{x}\right|\\ &\Rightarrow\left|\frac{(q_{x}-x_{a})}{\sqrt{(q_{x}-x_{a})^{2}+y_{a}^{2}+(z_{e}-z_{a})^{2}}}\right|=\left|\frac{(x_{n}-q_{x})}{\sqrt{(x_{n}-q_{x})^{2}+y_{n}^{2}+(z_{n}-z_{e})^{2}}}\right|\\ &\Rightarrow(q_{x}-x_{a})^{2}((x_{n}-q_{x})^{2}+y_{n}^{2}+(z_{n}-z_{e})^{2})=(x_{n}-q_{x})^{2}((q_{x}-x_{a})^{2}+y_{a}^{2}+(z_{e}-z_{a})^{2})\\ &\Rightarrow(\lambda(x_{1}-x_{2})+x_{2}-x_{a})^{2}((x_{n}-(\lambda(x_{1}-x_{2})+x_{2})))^{2}+y_{n}^{2}+(z_{n}-z_{e})^{2})\\ &\;\;\;\;\;\;\;=(x_{n}-(\lambda(x_{1}-x_{2})+x_{2}))^{2}(((\lambda(x_{1}-x_{2})+x_{2})-x_{a})^{2}+y_{a}^{2}+(z_{e}-z_{a})^{2})\\ &\Rightarrow a\lambda^{2}+b\lambda+c=0.\end{split}$ (21)
@@
$\begin{split}\bm{E}^{d}&=\begin{aligned} \left(E_{0x}\hat{\bm{\beta}}_{0x}^{\prime}D_{s}\right)\left(\hat{\beta}_{0x}\hat{x}+\hat{\beta}_{0y}\hat{y}+\hat{\beta}_{0z}\hat{z}\right)\\ \end{aligned}\frac{\mathrm{e}^{-jk(|s^{{}^{\prime}}|+|s|)}}{|s^{{}^{\prime}}|\sqrt{|s|}}.\end{split}$ (23)
Observe, by controlling the polarization, we have eliminated the hard diffraction coefficient from our electric field expression. Now, we calculate the ratio of the squared magnitude of this simplified electric field expression and then substitute the approximated soft diffraction coefficients derived in Lemma 2 and presented in Table IV. Note that $|s^{\prime}|$ and $|s|$ are the OPL and IPL defined in Corollary 1 respectively and are of similar length according to Assumption 2. The power ratio result follows from this.
@@
We set up a model of a brick building in a realistic wireless RayTracing software [42] as shown in Fig.  5. The UAV transmitter was placed 10m away from the exterior of the building. Further, we consider candidate receiver locations spread across five building floors, where each floor is located above the UAV transmitter height. Therefore, this would represent NLOS conditions for most of the locations on these floors. In line with the observation in Bas et al. [32], for each NLOS receiver location, we assume the direct path is sufficiently attenuated, and all the MPCs are generated due to interactions with the windows. We also assume that the windows are made of plain glass, so the attenuation offered to the impinging wireless signal is negligible [43]. The dimensions of each building floor are $20m\times 20m$, and we consider $N_{rx}=8000$ uniformly sampled receiver locations across the five floors of the building.
Figure 5: RayTracing setup using Remcom’s Wireless Insight [42] with the different MPC groups labeled on an AoA vs path length plot for an RX location on the $5^{th}$ floor of the building near the back.
For an arbitrary Rx location, the path followed by a particular MPC can be represented as a string such as ‘Tx-X-X-$\cdots$-X-Rx’. Since the MPC begins at the transmitter and terminates at the receiver, all possible strings start with a ‘Tx’ and terminate with an ‘Rx’. The character ‘X’ in the string denotes either ‘R’ or ‘D’ representing ‘Reflection’ or ‘Diffraction’. By reflection, we mean specular reflection from flat surfaces following Snell’s laws, whereas diffraction is from edges and follows the diffraction law in eq. (2). The number of characters ‘X’ between the ‘Tx’ and ‘Rx’ represents the number of interactions, and the left to right order of the characters denotes the sequence of interactions. For our model in Fig. 5, reflections happen at flat surfaces like ceilings, walls, and floors, whereas diffraction happens at edges like the window edges.
@@
String Tx-Rx Tx-R-Rx Tx-D-Rx Tx-R-R-$\cdots$-Rx or Tx-D-R-$\cdots$-Rx
$P_{e}$ $9.77\%$ $14.41\%$ $93.09\%$ $100\%$
$P_{fap}$ $9.77\%$ $0.23\%$ $82.30\%$ $7.67\%$
For each Rx location, we have many MPCs formed, and we group the MPCs in four unique groups: (a) MPC-1, (b) MPC-2, (c) MPC-3, and (d) MPC-4 based on the propagation mechanism and number of interactions associated with the group. Table VI shows the four possible MPC groups, with the string in the second row representing the corresponding propagation mechanism. Now we evaluate two metrics across all receiver locations (a) Probability of existence: $P_{e}$ and (b) Probability of first arrival path: $P_{fap}$ for the four MPC groups. Mathematically we can express $P_{e}$ and $P_{fap}$ as
$\begin{split}P_{e}=\frac{N_{e}\times 100}{N_{rx}},\;\;\;P_{fap}=\frac{N_{fap}\times 100}{N_{rx}}.\end{split}$ (24)
Here, $N_{e}$ is the number of Rx locations where a particular MPC group exists, and $N_{fap}$ is the number of Rx locations where the first arriving path belongs to a particular MPC Group. In Fig. 5, we also plot an AoA vs normalized path length plot for an Rx location at the back of the building. The normalized path length is calculated by subtracting the path length of the first arriving path. Hence, the length of the first arriving path is $0m$. The AoA is defined as the elevation angle from the Z-axis. We present three key observations about the different MPC groups based on Table VI and the AoA vs Path length plot.
@@
Remark.
From the associated string in Table VI, MPC-1 contains MPCs that follow the direct Euclidean path (LOS path) between the transmitter and receiver, whereas MPC-2 contains MPCs that undergo a single bounce satisfying Snell’s laws before reaching the receiver. The low probability of existence can be attributed to the fact that both these MPC groups only exist for Rx locations close to the windows.
@@
For MPC-3, we have a high probability of existence - $P_{e}=93.09\%$ and also a high probability that the first arriving path belongs to MPC-3 - $P_{fap}=82.3\%$. Lastly, the elevation AoA for MPC-3 is approximately $90$°.
Remark.
The raytracing simulation results show two diffraction paths being generated due to the upper and lower edges of the window on the Rx floor as predicted from the simplified building model presented in Section III-A. Further, from the path length vs AoA plot in Fig. 5, observe that the elevation AoA for the diffraction MPCs is 90°. This is evidence for proposition 1 where the diffraction MPCs arrive at the receiver parallel to the ground. For Rx locations occluded by the vertical pillar between windows, we would have no edge diffraction; instead, ‘Tx-D-Rx’ paths would be as a result of corner diffraction, which is not modeled in Wireless Insight, therefore $P_{e}\neq 100\%$. The first arriving path also happens to be contained in MPC-3 with the highest probability since $P_{fap}=82.3\%$. This is expected since the diffraction field follows the shortest path and was the basis for deriving the path length for the diffraction MPCs using Fermat’s principle of least time in Lemma 1.
@@
For MPC-4, the MPCs contained in this have varying path lengths and AoA.
Remark.
Since these MPCs undergo several interactions with the window and reflections from the insides of the building, we see varying path lengths and AoAs. Note that even if we were to account for corner diffractions, some MPCs with multiple interactions might be the shortest paths for a select few Rx locations. However, numerical results from the raytracing simulation show these Rx locations are in the minority.
@@
The entries in the Jacobian matrix $J$ in eq. (20) can be derived by using the chain rule on the path length $p_{j}(\bm{\alpha})$ between the $j^{th}$ anchor to the node at $\bm{\alpha}$ given by eq. (12). For each anchor, we obtain the x-coordinate $q_{x}$ of the diffraction point $Q_{e}$ on the diffracting edge by the expressions in Lemma 1 with the coefficients of the quadratic $a,b,c$ defined in eq. (4) for each anchor with $z_{e}=z_{n}+0.5w$. For notational simplicity, drop the subscript $j$, and argument $\bm{\alpha}$ in the expression for the entries in the Jacobian. Hence, the path length is denoted as $p$ and is defined for the anchor at $[x_{a},y_{a},z_{a}]^{T}$, the node at $[x_{n},y_{n},z_{n}]^{T}$ and the window size is $w$. The final expression is below.
$\begin{split}\frac{\partial p}{\partial x_{n}}&=\frac{(q_{x}-x_{a})\frac{\partial q_{x}}{\partial x_{n}}}{\sqrt{(q_{x}-x_{a})^{2}+(y_{a})^{2}+(z_{a}-z_{n}-0.5w)^{2}}}\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\frac{(x_{n}-q_{x})(1-\frac{\partial q_{x}}{\partial x_{n}})}{\sqrt{(x_{n}-q_{x})^{2}+(y_{n})^{2}+(0.5w)^{2}}},\\ \frac{\partial p}{\partial y_{n}}&=\frac{(q_{x}-x_{a})\frac{\partial q_{x}}{\partial y_{n}}}{\sqrt{(q_{x}-x_{a})^{2}+(y_{a})^{2}+(z_{a}-z_{n}-0.5w)^{2}}}\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\frac{(q_{x}-x_{n})\frac{\partial q_{x}}{\partial y_{n}}+y_{n}}{\sqrt{(x_{n}-q_{x})^{2}+(y_{n})^{2}+(0.5w)^{2}}},\\ \frac{\partial p}{\partial z_{n}}&=\frac{(q_{x}-x_{a})\frac{\partial q_{x}}{\partial z_{n}}}{\sqrt{(q_{x}-x_{a})^{2}+(y_{a})^{2}+(z_{a}-z_{n}-0.5w)^{2}}}\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\frac{(q_{x}-x_{n})\frac{\partial q_{x}}{\partial z_{n}}}{\sqrt{(x_{n}-q_{x})^{2}+(y_{n})^{2}+(0.5w)^{2}}},\\ \frac{\partial q_{x}}{\partial x_{n}}&=\frac{(x_{1}-x_{2})}{2a}\left[\frac{\partial a}{\partial x_{n}}\left(\frac{(b\mp\sqrt{b^{2}-4ac}}{a}\right)\right.\\ &\left.\;\;\;\;\;\;\;\;\;\;\;\;+\left(-\frac{\partial b}{\partial x_{n}}\pm\frac{b\frac{\partial b}{\partial x_{n}}-2c\frac{\partial a}{\partial x_{n}}-2a\frac{\partial c}{\partial x_{n}}}{\sqrt{b^{2}-4ac}}\right)\right],\\ \end{split}$
$\begin{split}\frac{\partial q_{x}}{\partial y_{n}}&=\frac{(x_{1}-x_{2})}{2a}\left[\frac{\partial a}{\partial y_{n}}\left(\frac{(b\mp\sqrt{b^{2}-4ac}}{a}\right)\right.\\ &\left.\;\;\;\;\;\;\;\;\;\;\;\;+\left(-\frac{\partial b}{\partial y_{n}}\pm\frac{b\frac{\partial b}{\partial y_{n}}-2c\frac{\partial a}{\partial y_{n}}-2a\frac{\partial c}{\partial y_{n}}}{\sqrt{b^{2}-4ac}}\right)\right],\\ \frac{\partial q_{x}}{\partial z_{n}}&=\frac{(x_{1}-x_{2})}{2a}\left[\frac{\partial a}{\partial z_{n}}\left(\frac{(b\mp\sqrt{b^{2}-4ac}}{a}\right)\right.\\ &\left.\;\;\;\;\;\;\;\;\;\;\;\;+\left(-\frac{\partial b}{\partial z_{n}}\pm\frac{b\frac{\partial b}{\partial z_{n}}-2c\frac{\partial a}{\partial z_{n}}-2a\frac{\partial c}{\partial z_{n}}}{\sqrt{b^{2}-4ac}}\right)\right],\\ \frac{\partial a}{\partial x_{n}}&=0,\;\frac{\partial a}{\partial y_{n}}=2y_{n}(x_{1}-x_{2})^{2},\frac{\partial a}{\partial z_{n}}=2(x_{1}-x_{2})^{2}(z_{n}+\frac{w}{2}),\\ \frac{\partial b}{\partial x_{n}}&=2(x_{1}-x_{2})\left((z_{1}-z_{a})^{2}+y_{a}^{2}\right),\\ \frac{\partial b}{\partial y_{n}}&=4y_{n}(x_{1}-x_{2})(x_{2}-x_{a}),\\ \frac{\partial b}{\partial z_{n}}&=-4(x_{1}-x_{2})(x_{2}-x_{n})(z_{n}+\frac{w}{2}-z_{a}),\\ \frac{\partial c}{\partial x_{n}}&=2(x_{2}-x_{n})\left((z_{1}-z_{a})^{2}+y_{a}^{2}\right),\\ \frac{\partial c}{\partial y_{n}}&=2y_{n}(x_{2}-x_{a})^{2},\frac{\partial c}{\partial z_{n}}=-2(x_{2}-x_{n})^{2}(z_{n}+\frac{w}{2}-z_{a}).\end{split}$ (25)

### v2

@@ VI Conclusion
This paper explores positioning in NLoS scenarios, focusing on diffraction as a signal propagation mechanism. Leveraging the Geometrical Theory of Diffraction (GTD), we propose a novel path model for NLoS scenarios where propagation paths are formed through interactions with edges. Based on this model, we develop a positioning technique for a public safety scenario addressing the need for improved 3D and z-axis position estimates where the O2I signal propagation is primarily due to diffraction from window edges through diffraction MPCs. Theoretical analysis based on GTD presents several insights about these diffraction MPCs, such as the presence of two diffraction MPCs resulting from diffraction from the upper and lower edges of the window located on the node building floor. In addition, these MPCs for diffraction arrive in parallel horizontal plane with AoA $\approx 90$°. Additionally, by changing the transmit side electric field polarization we can control the relative power in the two diffraction MPCs. In the second part of the paper, we establish the necessary and sufficient conditions for estimating the unknown 3D node position inside the building by using Fisher information analysis, and also derive the CRLB for 3D positioning accuracy for our proposed positioning technique. Further, we derive a non-linear least squares algorithm based on our proposed diffraction path model (D-NLS). We also present a realistic raytracing based study that analyses the different MPCs generated for our positioning scenario using statistical metrics. The generated MPCs are also used to generate numerical positioning results showing an improvement in positioning performance for our proposed positioning technique compared to the prior art.
@@
Consider Fig. 1(a); our goal is to derive the geometrical path length $\bm{X_{a}Q_{e}\alpha}$. We start by calculating the coordinates of the diffraction point $\bm{Q_{e}}$. For this, we substitute the incident ray unit vector $\bm{\hat{s^{\prime}}}$, edge unit vector $\bm{\hat{e}}$ and diffracted ray unit vector $\bm{\hat{s}}$ in terms of the various coordinates shown in Fig. 1(a) in the diffraction law eq. (2). Next, we substitute the parametric coordinates of the diffraction point $\bm{Q_{e}}$ in terms of the parameter $\lambda$ and the endpoints of the diffracting edge $\bm{X_{1}}$ and $\bm{X_{2}}$. After some algebraic manipulations, we obtain a quadratic equation for the parameter $\lambda$. All these steps are shown in eq. (24).
$\begin{split}&\left|\frac{(q_{x}-x_{a})\bm{\hat{x}}-y_{a}\bm{\hat{y}}+(z_{e}-z_{a})\bm{\hat{z}}}{\sqrt{(q_{x}-x_{a})^{2}+y_{a}^{2}+(z_{e}-z_{a})^{2}}}\cdot\bm{\hat{x}}\right|=\left|\frac{(x_{n}-q_{x})\bm{\hat{x}}+y_{n}\bm{\hat{y}}+(z_{n}-z_{e})\bm{\hat{z}}}{\sqrt{(x_{n}-q_{x})^{2}+y_{n}^{2}+(z_{n}-z_{e})^{2}}}\cdot\bm{\hat{x}}\right|\\ &\Rightarrow\left|\frac{(q_{x}-x_{a})}{\sqrt{(q_{x}-x_{a})^{2}+y_{a}^{2}+(z_{e}-z_{a})^{2}}}\right|=\left|\frac{(x_{n}-q_{x})}{\sqrt{(x_{n}-q_{x})^{2}+y_{n}^{2}+(z_{n}-z_{e})^{2}}}\right|\\ &\Rightarrow(q_{x}-x_{a})^{2}((x_{n}-q_{x})^{2}+y_{n}^{2}+(z_{n}-z_{e})^{2})=(x_{n}-q_{x})^{2}((q_{x}-x_{a})^{2}+y_{a}^{2}+(z_{e}-z_{a})^{2})\\ &\Rightarrow(\lambda(x_{1}-x_{2})+x_{2}-x_{a})^{2}((x_{n}-(\lambda(x_{1}-x_{2})+x_{2})))^{2}+y_{n}^{2}+(z_{n}-z_{e})^{2})\\ &\;\;\;\;\;\;\;=(x_{n}-(\lambda(x_{1}-x_{2})+x_{2}))^{2}(((\lambda(x_{1}-x_{2})+x_{2})-x_{a})^{2}+y_{a}^{2}+(z_{e}-z_{a})^{2})\\ &\Rightarrow a\lambda^{2}+b\lambda+c=0.\end{split}$ (24)
@@
$\begin{split}\bm{E}^{d}&=\begin{aligned} \left(E_{0x}\hat{\beta}_{0x}^{\prime}D_{s}\right)\left(\hat{\beta}_{0x}\bm{\hat{x}}+\hat{\beta}_{0y}\bm{\hat{y}}+\hat{\beta}_{0z}\bm{\hat{z}}\right)\\ \end{aligned}\frac{\mathrm{e}^{-jk(|\bm{s^{{}^{\prime}}}|+|\bm{s}|)}}{|\bm{s^{{}^{\prime}}}|\sqrt{|\bm{s}|}}.\end{split}$ (26)
Observe, by controlling the polarization, we have eliminated the hard diffraction coefficient from our electric field expression. Now, we calculate the ratio of the squared magnitude of this simplified electric field expression and then substitute the approximated soft diffraction coefficients derived in Lemma 2 and presented in Table IV. Note that $|\bm{s^{\prime}}|$ and $|\bm{s}|$ are the OPL and IPL defined in Corollary 1 respectively and are of similar length according to Assumption 2. The power ratio result follows from this.
@@
We set up a model of a building in a realistic wireless raytracing software [31] as shown in Fig.  6. Further, we consider candidate receiver (Rx) locations spread across five building floors, where each floor is located above the UAV anchor which is equipped with a radio transmitter (Tx). Note, for majority of the candidate Rx locations across floors this would represent NLoS conditions. According to the observation in Bas et al. [38], for each NLoS Rx location, we assume that the direct path is sufficiently attenuated and that all MPCs are generated due to interactions with the windows. We also assume that the windows are made of plain glass, so the attenuation offered to the impinging wireless signal is negligible [28].
Figure 6: Raytracing setup using Remcom’s Wireless Insight [31]. We create a building with floors, ceilings and exterior made out of ITU 28 GHz Concrete and Drywall partitions in the interiors. For numerical results we consider 17485 candidate receiver locations located across five building floors which are NLoS to all four UAV-anchors situated outside the building.
For an arbitrary Rx location, the path followed by a particular MPC can be represented as a string such as ‘Tx-X-X-$\cdots$-X-Rx’. Since the MPC begins at the transmitter and terminates at the receiver, all possible strings start with a ‘Tx’ and terminate with an ‘Rx’. The character ‘X’ in the string can be either ‘R’ denoting ‘Reflection’, ‘D’ denoting ‘Diffraction’ or ‘T’ denoting ‘Transmission’. By reflection, we mean reflection from flat smooth surfaces following Snell’s laws, transmission does not change the propagation direction and only attenuates the signal whereas diffraction is from edges and follows the diffraction law as in eq. (2). The number of characters ‘X’ between the ‘Tx’ and ‘Rx’ represents the number of interactions, and the left to right order of the characters denotes the sequence of interactions.
Since in our 3D positioning technique we use the first arrival path principle, we define the First Arrival Path between a fixed Tx and Rx as
Definition 3.
First Arrival Path: The particular MPC with the smallest ToF $\tau_{min}$ and SNR above a threshold $T$ among all existing MPCs at that location.
For the SNR calculation, we assume an Rx processing gain of 20 dB which is added to the signal power for each MPC extracted from the ray tracing simulation. Assuming a signal bandwidth of $B=400$ MHz and system temperature of $T=300$K, we obtain the noise floor as $N_{0}=KTB=-87$dBm. The threshold changes according to every pair of Rx-Tx according to $T=P_{max}-15$ where $P_{max}$ is the maximum SNR across all existing MPCs at a particular Rx location for a given Tx. Now we conduct statistical analysis on the MPCs generated by the raytracing simulation across all UAV transmitters and receivers. First, we look at the elevation AoA and SNR for the first arrival path.
(a) Elevation AoA Boxplot.png
(b) SNR Boxplot.png
Figure 7: First arriving path statistics across Rx locations. We can see that the elevation AoA is always within six degrees of $90$° and that the SNR is typically $>5$dB.
@@
Prop. Mech. Tx-Rx Tx-R-Rx Tx-D-Rx Tx-D-T-Rx Tx-D-T-T-Rx Tx-R-R-$\cdots$-Rx Tx-D-R-$\cdots$-Rx
$P_{e}$ $3.47\%$ $9.03\%$ $91.52\%$ $100\%$
$P_{fap}$ $3.47\%$ $1.60\%$ $84.79\%$ $10.12\%$
For a fixed Rx location and fixed Tx (UAV anchor), we can group the MPCs in four unique groups: (a) MPC-1, (b) MPC-2, (c) MPC-3, and (d) MPC-4 based on the propagation mechanism and number of interactions associated with the group. Table VI shows the four possible MPC groups, the string in the second row representing the corresponding propagation mechanism. Now we evaluate two metrics across all receivers and UAV anchors (a) Probability of existence: $P_{e}$ of the different MPC groups and (b) Probability that the first arrival path belongs to each MPC group: $P_{fap}$. Mathematically we can express $P_{e}$ and $P_{fap}$ as
$\begin{split}P_{e}=\sum_{tx}\frac{N_{e,tx}\times 100}{N_{rx}},\;\;\;P_{fap}=\sum_{tx}\frac{N_{fap,tx}\times 100}{N_{rx}}.\end{split}$ (27)
Here, $N_{e,tx}$ is the number of Rx locations for each transmitter (UAV anchor) where a particular MPC group exists, and $N_{fap,tx}$ is the number of Rx locations for each transmitter where the first arriving path belongs to a particular MPC Group. We present three key observations on the different MPC groups based on Table VI and Fig. 7.
@@
From the associated string in Table VI, MPC-1 contains MPCs that follow the direct Euclidean path (LoS path) between the transmitter and receiver, whereas MPC-2 contains MPCs that undergo a single bounce satisfying Snell’s laws before reaching the receiver. The low probability of existence can be attributed to the fact that both these MPC groups exist only for Rx locations close to the windows.
@@
For MPC-3 we have a high we have a high probability of existence $P_{e}$ and high probability that they are also first arriving paths since $P_{fap}=84.79\%$.
The results of the ray tracing simulation indicate that two diffraction paths are generated due to the upper and lower edges of the window on the receiver floor, consistent with the simplified building model described in Section III-A. The first arriving path belongs to MPC-3 with the highest probability ($P_{fap}=84.79\%$). This outcome aligns with expectations, as the diffraction field follows the shortest path, a principle used to derive the path length for diffraction MPCs based on Fermat’s principle of least time in Lemma 1. Furthermore, in Fig. 7, it can be observed that the first arriving path most likely corresponds to MPC-3, with an elevation AoA close to 90°. This observation supports Proposition 1, which states that diffraction MPCs arrive at the receiver parallel to the ground.
@@
The first arriving path principle rarely contains MPCs belonging to MPC-4.
Since these MPCs belonging to this group undergo several interactions with the window and reflections from the insides of the building, we see varying path lengths.
@@
The entries in the Jacobian matrix $J$ in eq. (20) can be derived by using the chain rule on the path length $p_{j}(\bm{\alpha})$ between the $j^{th}$ anchor to the node at $\bm{\alpha}$ given by eq. (12). For each anchor, we obtain the x-coordinate $q_{x}$ of the diffraction point $Q_{e}$ on the diffracting edge by the expressions in Lemma 1 with the coefficients of the quadratic equation $a,b,c$ defined in eq. (4) for each anchor with $z_{e}=z_{n}+0.5w$. For notational simplicity, drop the subscript $j$, and argument $\bm{\alpha}$ in the expression for the entries in the Jacobian. Hence, the path length is denoted as $p$ and is defined for the anchor at $\bm{X_{a}}=[x_{a},y_{a},z_{a}]^{T}$, the node at $\bm{\alpha}=[x_{n},y_{n},z_{n}]^{T}$ and the window size is $w$. The final expression is below.
$\begin{split}\frac{\partial p}{\partial x_{n}}&=\frac{(q_{x}-x_{a})\frac{\partial q_{x}}{\partial x_{n}}}{\sqrt{(q_{x}-x_{a})^{2}+(y_{a})^{2}+(z_{a}-z_{n}-0.5w)^{2}}}\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\frac{(x_{n}-q_{x})(1-\frac{\partial q_{x}}{\partial x_{n}})}{\sqrt{(x_{n}-q_{x})^{2}+(y_{n})^{2}+(0.5w)^{2}}},\\ \frac{\partial p}{\partial y_{n}}&=\frac{(q_{x}-x_{a})\frac{\partial q_{x}}{\partial y_{n}}}{\sqrt{(q_{x}-x_{a})^{2}+(y_{a})^{2}+(z_{a}-z_{n}-0.5w)^{2}}}\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\frac{(q_{x}-x_{n})\frac{\partial q_{x}}{\partial y_{n}}+y_{n}}{\sqrt{(x_{n}-q_{x})^{2}+(y_{n})^{2}+(0.5w)^{2}}},\\ \frac{\partial p}{\partial z_{n}}&=\frac{(q_{x}-x_{a})\frac{\partial q_{x}}{\partial z_{n}}}{\sqrt{(q_{x}-x_{a})^{2}+(y_{a})^{2}+(z_{a}-z_{n}-0.5w)^{2}}}\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\frac{(q_{x}-x_{n})\frac{\partial q_{x}}{\partial z_{n}}}{\sqrt{(x_{n}-q_{x})^{2}+(y_{n})^{2}+(0.5w)^{2}}},\\ \frac{\partial q_{x}}{\partial x_{n}}&=\frac{(x_{1}-x_{2})}{2a}\left[\frac{\partial a}{\partial x_{n}}\left(\frac{(b\mp\sqrt{b^{2}-4ac}}{a}\right)\right.\\ &\left.\;\;\;\;\;\;\;\;\;\;\;\;+\left(-\frac{\partial b}{\partial x_{n}}\pm\frac{b\frac{\partial b}{\partial x_{n}}-2c\frac{\partial a}{\partial x_{n}}-2a\frac{\partial c}{\partial x_{n}}}{\sqrt{b^{2}-4ac}}\right)\right],\\ \frac{\partial q_{x}}{\partial y_{n}}&=\frac{(x_{1}-x_{2})}{2a}\left[\frac{\partial a}{\partial y_{n}}\left(\frac{(b\mp\sqrt{b^{2}-4ac}}{a}\right)\right.\\ &\left.\;\;\;\;\;\;\;\;\;\;\;\;+\left(-\frac{\partial b}{\partial y_{n}}\pm\frac{b\frac{\partial b}{\partial y_{n}}-2c\frac{\partial a}{\partial y_{n}}-2a\frac{\partial c}{\partial y_{n}}}{\sqrt{b^{2}-4ac}}\right)\right],\\ \frac{\partial q_{x}}{\partial z_{n}}&=\frac{(x_{1}-x_{2})}{2a}\left[\frac{\partial a}{\partial z_{n}}\left(\frac{(b\mp\sqrt{b^{2}-4ac}}{a}\right)\right.\\ &\left.\;\;\;\;\;\;\;\;\;\;\;\;+\left(-\frac{\partial b}{\partial z_{n}}\pm\frac{b\frac{\partial b}{\partial z_{n}}-2c\frac{\partial a}{\partial z_{n}}-2a\frac{\partial c}{\partial z_{n}}}{\sqrt{b^{2}-4ac}}\right)\right],\\ \end{split}$
$\begin{split}\frac{\partial a}{\partial x_{n}}&=0,\;\frac{\partial a}{\partial y_{n}}=2y_{n}(x_{1}-x_{2})^{2},\frac{\partial a}{\partial z_{n}}=2(x_{1}-x_{2})^{2}(z_{n}+\frac{w}{2}),\\ \frac{\partial b}{\partial x_{n}}&=2(x_{1}-x_{2})\left((z_{1}-z_{a})^{2}+y_{a}^{2}\right),\\ \frac{\partial b}{\partial y_{n}}&=4y_{n}(x_{1}-x_{2})(x_{2}-x_{a}),\\ \frac{\partial b}{\partial z_{n}}&=-4(x_{1}-x_{2})(x_{2}-x_{n})(z_{n}+\frac{w}{2}-z_{a}),\\ \frac{\partial c}{\partial x_{n}}&=2(x_{2}-x_{n})\left((z_{1}-z_{a})^{2}+y_{a}^{2}\right),\\ \frac{\partial c}{\partial y_{n}}&=2y_{n}(x_{2}-x_{a})^{2},\frac{\partial c}{\partial z_{n}}=-2(x_{2}-x_{n})^{2}(z_{n}+\frac{w}{2}-z_{a}).\end{split}$ (28)